# Generative Anchored Fields: Controlled Data Generation via Emergent Velocity Fields and Transport Algebra

### v1

@@ Abstract
We present Generative Anchored Fields (GAF), a generative model that learns independent endpoint predictors $J$ (noise) and $K$ (data) rather than a trajectory predictor. The velocity field v=KJv=K-J emerges from their time-conditioned disagreement. This factorization enables Transport Algebra: algebraic operation on learned $\{(J_{n},K_{n})\}_{n=1}^{N}$ heads for compositional control. With class-specific $K_{n}$ heads, GAF supports a rich family of directed transport maps between a shared base distribution and multiple modalities, enabling controllable interpolation, hybrid generation, and semantic morphing through vector arithmetic. We achieve strong sample quality (FID 7.5 on CelebA-HQ $64\times 64$) while uniquely providing compositional generation as an architectural primitive. We further demonstrate, GAF has lossless cyclic transport between its initial and final state with LPIPS=$0.0$. Code available at https://github.com/IDLabMedia/GAF
@@ 1 Introduction
Modern generative models [kingma_auto-encoding_2013, goodfellow_generative_2014, ho_denoising_2020] achieve remarkable sample quality but lack precise compositional control. Score-based Diffusion and flow matching models learn a single time-conditioned predictor that maps noise to data via a unified field [pmlr-v37-sohl-dickstein15, lipman_flow_2023]. Although models such as Stable Diffusion [Rombach_2022_CVPR, podell2024sdxl] and FLUX.1 [blackforestlabs2025flux] generate images with high fidelity, their monolithic architecture treats control as an external steering process, implemented through classifier-free guidance [ho2021classifierfree], prompt engineering, latent space editing [mokady2022null], or attention manipulation [hertz2022prompt], rather than an intrinsic property of the learned representation.
Figure 1: Generated samples demonstrating our transport algebra framework. (a) Selected samples from CelebA-HQ ($256\times 256$ px), demonstrating high-fidelity generation across diverse demographics and attributes. (b) Lossless cyclic transport across three domains: $v_{\text{cat}_{0}}\to v_{\text{dog}}\to v_{\text{wild}}\to v_{\text{cat}_{0}}$ with LPIPS=$0.0$. (c) Barycentric interpolation: $v_{i\to j\to k}=\alpha v_{i}+\beta v_{j}+\gamma v_{k}$, showing smooth multi-domain transitions in velocity space.
These limitations suggest a need to re-examine the core learning objectives. The trajectory-based paradigm focuses on learning step-by-step dynamics along a path [liu2022flow]. This raises a fundamental question: must we learn a step-by-step mechanism of the trajectory or is it sufficient to know only its origin and destination? The trajectory-based paradigm assumes the first option. We explore the second one.
@@
This factorization has a direct consequence: precise and controllable generation. Because GAF learns independent endpoint predictors $\{(J_{n},K_{n})\}_{n=1}^{N}$, particularly when using class-specific $K_{n}$ heads, it naturally enables compositional operations at inference time. One can switch between different $K$ heads mid-trajectory, interpolate between class manifolds with explicit geometric control, or arithmetically combine predictors to generate novel outputs. These operations are architectural primitives in GAF. We term this capability Transport Algebra: the ability to perform algebraic manipulations directly on the model’s learned components to achieve precise semantic control.
@@
  • Transport Algebra. GAF introduces transport algebra to generative modeling: the ability to perform arithmetic operations directly on learned endpoint predictors. This enables controllable test-time generation, such as interpolation between classes, switching manifolds mid-trajectory, or combining multiple classes for novel generation.
  • Emergent dynamics via endpoint-anchored learning. Instead of directly learning a score, noise or velocity field, GAF trains two time-conditioned endpoint predictors $J$ and $K$ that target a noise and data endpoint, respectively. The transport velocity field $v(\mathbf{x}_{t},t)=K-J$ emerges as their antisymmetric disagreement. This decouples the training objective (endpoint accuracy) from the sampler, which performs Ordinary Differential Equation (ODE) integration.
@@
  • Simple training and deterministic sampling. Training is a direct endpoint regression with simple residual and swap regularizers. Sampling is deterministic and guidance free, requiring only single-step Euler integration due to GAF’s self-correcting dynamics, unlike multi-step ODE solvers needed by other flow-based models..
@@ 2 Related Work
Modern DGM is largely dominated by methods that learn to reverse a data-to-noise trajectory, with key differences arising in the parametrization of this reverse process [song2021scorebased, albergo2023building, lipman_flow_2023]. These approaches can be broadly categorized into score-based, denoising and velocity-based formulations. Our work builds upon the velocity-based formalism of generative processes but introduces a novel endpoint regression objective for the underlying transport dynamics.
@@
Score-based generative models learn the score function (i.e., the gradient of the log-density $\nabla_{\mathbf{x}}\log p(\mathbf{x}_{t})$) of noise-perturbed data distribution [10.5555/3454287.3455354, 10.5555/3495724.3496767]. By training a noise conditional score network across multiple noise scales, these models can generate samples via annealed Langevin dynamics. This paradigm was later formalized through the lens of stochastic differential equations (SDE), which established the score function as the key component needed to reverse a continuous-time forward diffusion process [song2021scorebased]. This SDE framework provided a unifying mathematical language for a wide variety of diffusion models, but the core objective remained the estimation of the score.
@@
Concurrently, Denoising Diffusion Probabilistic Models (DDPMs) [pmlr-v37-sohl-dickstein15, ho_denoising_2020] achieved state-of-the-art sample quality by simplifying the transport objective to predict the noise $\epsilon$ added at each step of a discrete forward process. This noise-prediction ($\epsilon$-prediction) target proved highly stable and effective, becoming the default standard, such as in OpenAI’s DALL-E [dhariwal_diffusion_2021], Google’s Imagen [saharia_photorealistic_2022], and Midjourney [pmlr-v162-nichol22a]. Later, Denoising Diffusion Implicit Models (DDIM) [song2021denoising] introduced deterministic sampling for diffusion models to remove randomness during generation, allowing the model to skip steps and produce high-quality images. Subsequent improvements focused on optimization, the training recipe, and architecture, with new methods like cosine noise schedules, learned variances [nichol_improved_2021], architectural upgrades [dhariwal_diffusion_2021], and powerful conditioning mechanisms [ho2021classifierfree]. A pivotal development was classifier-free guidance (CFG) [ho2021classifierfree], which enabled high-fidelity conditional generation from a single model by extrapolating between conditional and unconditional outputs. While our work uses the powerful transformer backbone (DiT) [vaswani_attention_2017, dosovitskiy2021an, williampeebles2023] that pioneered in this domain, we depart from the standard practice of predicting either the score or the noise.
@@
Most recently, the field has increasingly shifted towards a deterministic, continuous-time perspective rooted in ODEs and optimal transport [villani2009optimal, lipman_flow_2023, liu2022flow, lipman2024flowmatchingguidecode]. These models learn a velocity field $v_{\theta}(\mathbf{x}_{t},t)$ that governs a probability flow ODE, transporting samples from a simple base distribution (such as a Gaussian distribution) to the data distribution. While mathematically related to the score, this velocity ($v$-prediction) parametrization was shown to have better numerical stability and scaling properties [salimans2022progressive, 10.5555/3600270.3602196]. The development of simulation-free training objectives, such as Flow Matching [lipman_flow_2023] and Rectified Flows [liu2022flow], has made this approach highly efficient. By directly regressing the network’s output onto a target field, often defining a straight path between noise and data, these methods learn linear trajectories, enabling high-quality generation in very few sampling steps, as demonstrated by models like FLUX.1 [blackforestlabs2025flux], Stable Diffusion 3 [esser_scaling_2024], and InstaFlow [ICLR2024_4dc37a7b].
All these approaches, whether predicting score, noise, or velocity, learn a single time-conditioned predictor. While effective for generation, this monolithic architecture limits compositional control: one cannot independently manipulate learned representations for different classes or combine them arithmetically. Conditional generation instead relies on guidance mechanisms like CFG or prompt engineering.
@@
Recent work has explored alternative control mechanism. For example, Compositional Visual Generation [liu2022compositional] defines two compositional operators (AND and NOT) for multi-concept generation, using a set of diffusion models to enabling logical operation over concepts. Moreover, ControlNet [10377881] enables spatial control in text-to-image generation by locking the original diffusion model while training another network for conditional control with extra inputs like edges or depth. Although these methods provide impressive control, they require either maintaining multiple models for composing their outputs or training additional conditioning networks. Both approaches treat control as an add-on mechanism rather than an intrinsic architectural property.
@@
To address this limitation, GAF introduces an algebraic factorization of the transport dynamics. Instead of learning a single, monolithic velocity field $v_{\theta}$, GAF learns two endpoint predictors: a twin $J$ that regresses towards the noise distribution and a twin $K$ that regresses towards the data manifold. The velocity field $v$ then emerges from their difference, v=KJv=K-J, enforced by a paired and time-antisymmetric loss that promotes consistency between the forward and backward (reverse in time) dynamics. This explicit factorization of the transport operators ($J$ and $K$) enables unique architectural choices, such as independent, class-specialized $K$-heads. Unlike external composition [liu2022compositional] or auxiliary networks [10377881], GAF’s compositional capabilities are intrinsic: they emerge from the factorized architecture. This design unlocks novel inference-time capabilities, allowing vector operations between class distributions, a process we now term Transport Algebra. This form of algebraic manipulation over $\{(J_{n},K_{n})\}_{n=1}^{N}$ to generate new samples with precise compositional control is not naturally afforded by standard score-, flow-, or noise-based conditional generative models.
@@ 3 Generative Anchored Fields (GAF)
3.1 The GAF Framework
We introduce the Generative Anchored Fields model. Let $\mathbf{z}_{x},\mathbf{z}_{y}\in\mathbb{R}^{d}$ be samples from a data distribution $\mathbf{z}_{x}\sim p_{\text{data}}(\mathbf{x})$ and a standard normal distribution $\mathbf{z}_{y}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, respectively. Additionally, let $\mathbf{c}\in{\mathbb{R}^{d}}$ be an optional conditioning vector (e.g., a class embedding).
@@ 3.1.1 Velocity Field and ODE
We generate new data samples by integrating this ODE from $t=0$ to $t=1$ along either a linear or cosine time grid, using a single-step Euler integration [chen_neural_2018, 10.5555/3600270.3602196].
@@ 3.1.2 GAF Training
To improve training stability, we also enforce the model’s boundary behavior. That is, we examine the prediction errors at the ideal boundary condition $J-\mathbf{z}_{y}$ and $K-\mathbf{z}_{x}$ by expanding them at the endpoints.
@@
The endpoint residual penalty ensures $J\rightarrow\mathbf{z}_{y}$ as $t\rightarrow 0$ and $K\rightarrow\mathbf{z}_{x}$ as $t\rightarrow 1$, but it does not fully determine the intermediate behavior of GAF on the bridge. Since we want GAF to behave consistently across the entire bridge, we need a mechanism to do that. One way to do that is by swapping the residuals $J_{\text{res}}$ and $K_{\text{res}}$ across the bridge and create antisymmetric relationship between them. We define the swap operator (function) acting on the residuals at $t$ and 1t1-t as:
@@
As we have discussed earlier, in GAF, we do not train the velocity field directly, as in rectified flow. Instead, the velocity simply emerges from the twins opposite movement. However, from (21) and (22), we can see that the emergent velocity field incorporates knowledge of the trajectory’s endpoint.
@@
The overall architecture of GAF and its data flow is summarized in Figure 2.
@@ 3.2.1 Main Components
    Sampler (single ODE solver). Sampling integrates the emergent field with a single ODE solver (e.g., Heun sampler). No auxiliary corrector network, guidance model, or second model is required. Adding an additional learned sampler is possible but would depart from minimal GAF design.
@@ 3.2.2 GAF is Naturally Modular
GAF is modal-agnostic. That is, to change the modality of GAF, we only need to change the trunk’s modality and the I/O. For example, to change the modality from image data to video data, you only need to change the trunk. The other components need no (or minimal) change.

### v2

@@ Abstract
We present Generative Anchored Fields (GAF), a generative model that learns independent endpoint predictors, $J$ (noise) and $K$ (data), from any point on a linear bridge. Unlike existing approaches that use a single trajectory or score predictor, GAF is trained to recover the bridge endpoints directly via coordinate learning. The velocity field v=KJv=K-J emerges from their time-conditioned disagreement. This factorization enables Transport Algebra: algebraic operations on multiple $J/K$ heads for compositional control. With class-specific $K_{n}$ heads, GAF defines directed transport maps between a shared base noise distribution and multiple data domains, allowing controllable interpolation, multi-class composition, and semantic editing. This is achieved either directly on the predicted data coordinates ($K$) using Iterative Endpoint Refinement (IER), a novel sampler that achieves high-quality generation in 585-8 steps, or on the emergent velocity field ($v$). We achieve strong sample quality (FID 7.51 on ImageNet $256\times 256$ and $7.27$ on CelebA-HQ $256\times 256$, without classifier-free guidance) while treating compositional generation as an architectural primitive. Code available at https://github.com/IDLabMedia/GAF.
@@ 1 Introduction
Modern generative models [16, 10, 13] achieve remarkable sample quality but lack precise compositional control, which is the ability to independently manipulate, interpolate, or combine learned class representations at inference time. Score-based diffusion and flow matching models learn a single time-conditioned predictor that maps noise to data via a unified field [38, 20]. Although models such as Stable Diffusion [32, 31] and FLUX.1 [3] generate images with high fidelity, their monolithic architecture treats control as an external steering process, implemented through classifier-free guidance [14], prompt engineering, latent space editing [26], or attention manipulation [11], rather than an intrinsic property of the learned representation.
Figure 1: Overview of GAF. (a) Samples from CelebA-HQ ($256\times 256$ px). (b) Multi-class composition with custom masks. (c) Barycentric interpolation: $v_{i\to j\to k}=\alpha v_{i}+\beta v_{j}+\gamma v_{k}$; corners are pure classes, interior points are weighted mixtures.
These limitations suggest a need to re-examine the core learning objectives. The trajectory-based paradigm focuses on learning step-by-step dynamics along a path [23]. This raises a fundamental question: must we learn a step-by-step mechanism of the trajectory or is it sufficient to know only its origin and destination? The trajectory-based paradigm assumes the first option. We explore the second one.
@@
This factorization of the transport operators ($J$ and $K$) has a direct consequence: precise and controllable generation. Because GAF learns independent endpoint predictors $\{(J_{n},K_{n})\}_{n=1}^{N}$, particularly when using class-specific $K_{n}$ heads, it naturally enables compositional operations at inference time. For example, one can switch between different $K$ heads mid-trajectory, interpolate between class manifolds with explicit geometric control, or arithmetically combine predictors to generate novel outputs. These operations are architectural primitives in GAF. We term this capability Transport Algebra: the ability to perform algebraic manipulations directly on the model’s learned components to achieve precise semantic control.
@@
  • Coordinate-based Generative Modeling. Instead of learning a score, noise or velocity field, GAF learns the explicit coordinates of the transport endpoints: $J$ (noise) and $K$ (data). The velocity field $v(\mathbf{x}_{t},t)=K-J$ emerges as their antisymmetric disagreement. This formulation decouples the training objective (endpoint accuracy) from the sampler, allowing generation via Ordinary Differential Equation (ODE) integration or Iterative Endpoint Refinement (IER).
  • Transport Algebra. GAF introduces transport algebra to generative modeling: the ability to perform arithmetic operations directly on learned endpoint predictors. This enables controllable test-time generation, such as interpolation between classes, switching manifolds mid-trajectory, or combining multiple classes directly on the $K$ heads for novel compositional generation.
  • Iterative Endpoint Refinement. GAF has a native non-ODE sampler which iteratively refines endpoint estimates on a linear bridge, converging in 585-8 steps compared to 8025080-250 steps for Euler integration. IER forward uses only $K$ to refine the data endpoint, while IER reverse uses only $J$ to refine the noise endpoint, demonstrating that both the noise head $J$ and the data head $K$ are equally important.
[…]
@@
  • Simple training and deterministic sampling. Training is a direct endpoint regression with simple residual and swap regularizers. Sampling is deterministic and guidance free, requiring small step Euler integration due to GAF’s self-correcting dynamics, unlike multi-step ODE solvers needed by other flow-based models or using IER.
@@ 2 Related Work
Modern Deep Generative Modeling (DGM) is largely dominated by methods that learn to reverse a data-to-noise trajectory, with key differences arising in the parametrization of this reverse process [43, 2, 20]. These approaches can be broadly categorized into score-based, denoising and velocity-based formulations. Our work departs from these formulations by introducing an endpoint regression objective for underlying transport dynamics.
@@
Score-based generative models learn the score function (i.e., the gradient of the log-density $\nabla_{\mathbf{x}}\log p(\mathbf{x}_{t})$) of noise-perturbed data distribution [41, 42]. By training a noise conditional score network across multiple noise scales, these models can generate samples via annealed Langevin dynamics. This paradigm was later formalized through the lens of stochastic differential equations (SDE), which established the score function as the key component needed to reverse a continuous-time forward diffusion process [43]. This SDE framework provided a unifying mathematical language for a wide variety of diffusion models, but the core objective remained the estimation of the score.
@@
Concurrently, Denoising Diffusion Probabilistic Models (DDPMs) [38, 13] achieved state-of-the-art sample quality by simplifying the transport objective to predict the noise $\epsilon$ added at each step of a discrete forward process. This noise-prediction ($\epsilon$-prediction) target proved highly stable and effective, becoming the default standard, such as in OpenAI’s DALL-E [7], Google’s Imagen [34], and Midjourney [28]. Later, Denoising Diffusion Implicit Models (DDIM) [39] introduced deterministic sampling for diffusion models to remove randomness during generation, allowing the model to skip steps and produce high-quality images. Subsequent improvements focused on optimization, the training recipe, and architecture, with new methods like cosine noise schedules, learned variances [27], architectural upgrades [7], and powerful conditioning mechanisms [14]. A pivotal development was classifier-free guidance (CFG) [14], which enabled high-fidelity conditional generation from a single model by extrapolating between conditional and unconditional outputs. While our work uses the powerful transformer backbone (DiT) [45, 8, 30] that pioneered in this domain, we depart from the standard practice of predicting either the score or the noise.
@@
Most recently, the field has increasingly shifted towards a deterministic, continuous-time perspective rooted in ODEs and optimal transport [46, 20, 23, 21]. These models learn a velocity field $v_{\theta}(\mathbf{x}_{t},t)$ that governs a probability flow ODE, transporting samples from a simple base distribution (such as a Gaussian distribution) to the data distribution. While mathematically related to the score, this velocity ($v$-prediction) parametrization was shown to have better numerical stability and scaling properties [36, 15]. The development of simulation-free training objectives, such as Flow Matching [20] and Rectified Flows [23], has made this approach highly efficient. Stochastic Interpolants [1] generalize this framework by constructing interpolating paths between arbitrary distributions, sharing GAF’s bridge formulation but learning a velocity field rather than endpoint coordinates. Consistency Models [40] target few-step generation by distilling a pretrained diffusion model into a single-step predictor; GAF’s IER achieves comparable step efficiency (585-8 steps) without a pretrained teacher or distillation stage. By directly regressing the network’s output onto a target field, often defining a straight path between noise and data, these methods learn linear trajectories, enabling high-quality generation in very few sampling steps, as demonstrated by models like FLUX.1 [3], Stable Diffusion 3 [9], and InstaFlow [24].
@@
All these approaches, whether predicting score, noise, or velocity, learn a single time-conditioned predictor. While effective for generation, this monolithic architecture limits compositional control: one cannot independently manipulate learned representations for different classes or combine them arithmetically. Conditional generation instead relies on guidance mechanisms like CFG or prompt engineering. Recent work has explored alternative control mechanism. For example, Compositional Visual Generation [22] defines two compositional operators (AND and NOT) for multi-concept generation, using a set of diffusion models to enabling logical operation over concepts. Moreover, ControlNet [47] enables spatial control in text-to-image generation by locking the original diffusion model while training another network for conditional control with extra inputs like edges or depth. Although these methods provide impressive control, they require either maintaining multiple models for composing their outputs or training additional conditioning networks. Both approaches treat control as an add-on mechanism rather than an intrinsic architectural property.
@@
To address this limitation, GAF introduces an algebraic factorization of the transport dynamics. Instead of learning a single, monolithic velocity field $v_{\theta}$, GAF learns two endpoint predictors: a twin $J$ that regresses towards the noise distribution and a twin $K$ that regresses towards the data manifold. The velocity field $v$ then emerges from their difference, v=KJv=K-J, enforced by a paired and time-antisymmetric loss that promotes consistency between the forward and backward (reverse in time) dynamics. This explicit factorization of the transport operators ($J$ and $K$) enables unique architectural choices, such as independent, class-specialized $K$-heads. Unlike external composition [22] or auxiliary networks [47], GAF’s compositional capabilities are intrinsic: they emerge from the factorized architecture. This design unlocks novel inference-time capabilities, allowing vector operations between class distributions, a process we term Transport Algebra. This form of algebraic manipulation over $\{(J_{n},K_{n})\}_{n=1}^{N}$ to generate new samples with precise compositional control directly is not naturally afforded by standard score-, flow-, or noise-based conditional generative models.
A summary of how GAF compares to dominant trajectory-based paradigms is provided in Table 1.
Table 1: A comparison of GAF with Score-Based Diffusion and Flow Matching models across key aspects.
Feature Score-Based Diffusion Flow Matching GAF
Foundation Stochastic (SDE) Deterministic (ODE) Deterministic (IER/ODE)
Primary Learning Target A Score Function $\nabla_{x}\log{p}_{\theta}(x)$ Velocity Field ($v$) Coordinate Learning: Endpoints via twins, $J\!\to\!\mathbf{z}_{y},\ K\!\to\!\mathbf{z}_{x}$
Nature of Dynamics Derived (Indirect) Prescribed (Direct) Direct (IER); Emergent ($v$)
Velocity’s Role Derived from the score function. The direct regression target. A byproduct of endpoints (KJK-J)
Model’s Task Denoising/Score estimation Velocity Regression Endpoint Regression
@@ 3 Generative Anchored Fields (GAF)
GAF is a generative model that learns independent endpoint predictors rather than a trajectory or velocity field. In Section 3.1, we define the bridge formulation, the twin predictors $J$ and $K$, and derive the training objective. In Section 3.2, we discuss the structural properties that follow from this factorization. In Section 3.3, we introduce Transport Algebra, the compositional operations enabled by independent endpoint heads.
3.1 The GAF Model
We introduce the Generative Anchored Fields model. Let $\mathbf{z}_{x},\mathbf{z}_{y}\in\mathbb{R}^{d}$ be samples from a data distribution $\mathbf{z}_{x}\sim p_{\text{data}}(\mathbf{x})$ and a standard normal distribution $\mathbf{z}_{y}\sim\mathcal{N}(0,I)$, respectively. Additionally, let $\mathbf{c}\in{\mathbb{R}^{d}}$ be an optional conditioning vector (e.g., a class embedding).
@@ 3.1.1 Emergent Transport Dynamics
We generate new data samples by either (i) iteratively refining the endpoints via the IER algorithm or (ii) integrating the velocity field ODE from $t=0$ to $t=1$ along a linear or cosine time grid, using Euler integration [4, 15] (see Section 4.4).
@@ 3.1.2 GAF Training
Residual penalty. To improve training stability, we also enforce the model’s boundary behavior. That is, we examine the prediction errors at the ideal boundary condition $J-\mathbf{z}_{y}$ and $K-\mathbf{z}_{x}$ by expanding them at the endpoints.
@@
Residual-antisymmetry penalty. The endpoint residual penalty ensures $J\rightarrow\mathbf{z}_{y}$ as $t\rightarrow 0$ and $K\rightarrow\mathbf{z}_{x}$ as $t\rightarrow 1$, but it does not fully determine the intermediate behavior of GAF on the bridge. Since we want GAF to behave consistently across the entire bridge, we need a mechanism to do that. One way to do that is by swapping the residuals $J_{\text{res}}$ and $K_{\text{res}}$ across the bridge and create antisymmetric relationship between them. We define the swap operator (function) acting on the residuals at $t$ and 1t1-t as:
@@
As we have discussed earlier, in GAF, we do not train the velocity field directly, as in rectified flow (see Section 5.7 for a comparison). Instead, the velocity simply emerges from the twins opposite movement. However, from (21) and (22), we can see that the emergent velocity field incorporates knowledge of the trajectory’s endpoint.
@@
The overall architecture of GAF and its data flow is summarized in Figure 2, illustrating how the training components (top) predict endpoints that drive sampling strategies (bottom) via IER or velocity integration.
@@ 3.2.1 Main Components
    Sampler. GAF samples natively via Iterative Endpoint Refinement (IER), which refines endpoint estimates direclty on the bridge withouth solving an ODE. Alternatively the emergent field v=KJv=K-J can be integrated with a single ODE solver (e.g., Euler or Heun). In both cases, no auxiliary corrector network, guidance model, or second model is required.
@@ 3.2.2 GAF is Naturally Modular
GAF is modal-agnostic. That is, to change the modality of GAF, we only need to change the trunk’s modality and the I/O. For example, to change the modality from image data to video data, you only need to change the trunk. The other GAF components need no (or minimal) change.